Please help on physics question
An aluminum bar is moved across conducting rails a
Answer #1

motional emf e = i*R = B*L*v


i isthe current = 0.5 A

R = 6 ohm

B = 2.5 T

L = 1.2 m

v is the required speed

then i*R = B*L*v


v = i*R/(B*L) = (0.5*6)/(2.5*1.2) = 1 m/s